if a number is added to its square, the result is 56. find the number.
Found 3 solutions by mananth, jsmallt9, nabla:
Answer by mananth(16949)

Let the number be x
the other number will be x^2
x+x^2=56
x^2+x-56=0
x*2+8x-7x-56=0
x(x+8)-7(x+8)=0
(x+8)(x-7)=0

x=-8 or 7
x=7 satisfies the condition

Answer by jsmallt9(3759)

A solution provided by another tutor finds that
x = -8 or x = 7
And for some reason he/she suggests that x = -8 does not fit the problem. Both numbers are correct, not just x = 7.

Answer by nabla(475)

The question asks for a single number i.e. an element of the solution set. Both -8 and 7 fit this criteria, so either in fact can be offered as an answer that follows from a proper understanding of the question.